OpenStudy (anonymous):

There are 12 players on the basketball roster. 9 of the players are in the 8th grade. What percent of the team is not in the 8th grade?

OpenStudy (anonymous):

new question: There ae 12 players on the basketball roster. 9 of the players are in the 8th grade. what percent of the team is not in the 8th grade?

OpenStudy (anonymous):

I would suggest putting into fraction/simplifying it then turning into demals, from there easily into percentage

OpenStudy (anonymous):

decimals*

OpenStudy (anonymous):

can you do that or do you need more assistance into this matter

OpenStudy (anonymous):

sorry about puntuation mistakes, am in a hurry

OpenStudy (anonymous):

There are 12 players in whole team and 9 in G8 so 9over 12\[\left(\begin{matrix}9 \\ 12\end{matrix}\right)\] now we simplify these both numbers go into 3 so you get \[\left(\begin{matrix}3 \\ 4\end{matrix}\right)\] now into decimals so you get 0.75 which turns into 75%

OpenStudy (anonymous):

a store is having a 10% off sale. it gives a additional 20% off if the item is still above 100 dollars after the first discount. Jacket costs 124.99, if you get the second discount how many dollars have you saved in total?

OpenStudy (anonymous):

Are you sure that is right?

OpenStudy (anonymous):

a store is having a 10% off sale. it gives a additional 20% off if the item is still above 100 dollars after the first discount. Jacket costs 124.99, if you get the second discount how many dollars have you saved in total?

OpenStudy (anonymous):

hmm lol not really

OpenStudy (anonymous):

One of us misread the question, just saying... but that is a good way to solve it.

OpenStudy (anonymous):

i was rechecking now

OpenStudy (anonymous):

oh my bad yeh i worked how many are in in the 8th grade if im correct?

OpenStudy (anonymous):

yep, so what do you think it should be, Call?

OpenStudy (anonymous):

25%

OpenStudy (anonymous):

if 75% are in the 8th grade, what percent are not?

OpenStudy (anonymous):

two ways minus 75% from 100% or you could do it again with \[\left(\begin{matrix}3 \\ 12\end{matrix}\right)\]

OpenStudy (anonymous):

Call as your not involved in the discussion or part of the learning process and just waiting for the answers i refuse to help you.